OpenStudy (anonymous):

Tania went to America She changed £325 into dollars The exchange rate was £1 = $1.68 Change £325 into dollars

OpenStudy (anonymous):

with working out

OpenStudy (anonymous):

£1 = $1.68 so £325=325*1.68= $ 546

OpenStudy (anonymous):

325*(1.68/1) = values in $

OpenStudy (anonymous):

alright the second part is: When she came home she changed $117 into pounds The new exchange rate was £1 = $1.50 Change $117 into pounds

OpenStudy (anonymous):

it's the same way - 117*(1/1.5) = value in pounds The main concept here is using the units given and ratio (exchange rate)

OpenStudy (anonymous):

ok then thanks :D